OpenStudy (dbwong):

what is z? z +ln z = 1+e

OpenStudy (dbwong):

wolfram said you cant solve it algebraically, and i dont see a way to anyways.....

OpenStudy (itz_sid):

I think i might know... \[e(z+\ln z=1+e)\] \[ze + z = e + e^2\] \[z(e+1) = e+e^2\] \[z = \frac{ e+e^2 }{ e+1 }\] Can someone verify me? I took a shot at it.

OpenStudy (dbwong):

but eln(z) is not equal to z, your mistaking e^ln(z)

OpenStudy (dbwong):

e^ln(4) = 4 BUT if e*ln(4) does not equal 4

OpenStudy (irishboy123):

if \(z = e\), then \(\ln z = 1\) "a" solution

OpenStudy (tkhunny):

There is NOT a closed-form, expression that represents an algebraic solution. There is a numerical solution, as suggested by IrishBoy123. z + ln(z) is monotonic increasing, so we will not be finding another solution.

OpenStudy (loser66):

@iTz_Sid If you multiply e to the equation, then the second guy is elnz = ln z^e, not = z

OpenStudy (mathmale):

Newton's Method would help you to find one or more roots of z +ln z = 1+e. If you know this Method, start out by forming the new function f(z)=z+ln z-1-e.